OpenStudy (anonymous):

solve this in simplest form 14x2y3. <-- the 2 and 3 are exponents. over 35xy2 <-- exponent too. so 14x2y3/35xy2

OpenStudy (anonymous):

14x^(2)y^(3)/35(x)y^(2) 1/35x(y)^(2) = x^(-1)y^(-2)/35

OpenStudy (anonymous):

remember the rule that 1/x = x^(-1) and when multiplying variables with exponents you add them

OpenStudy (anonymous):

so bring the variables to the numerator 14x^(2)y^(3)x^(-1)y^(-2)/35

OpenStudy (anonymous):

now add the exponents to simplify, you do the next step im lazy